# Detecting Invariant Manifolds in ReLU-Based RNNs

### v1

@@ 3.3 Locating un-/stable manifolds
Input: $N_{max}$: Maximum number of iterations
    $P:$ periodic point
    $\sigma$: Index indicating whether the stable or unstable manifold is to be computed
    $E^{\sigma}$: Stable/unstable eigenvectors
    $\lambda^{\sigma}$: Stable/unstable eigenvalues
Output: $\{S^{\sigma}_{k},Q^{\sigma}_{k}\}_{k=0}^{k_{max}}$: Sampled points and expression for the manifold in each subregion
1:k=0 $\triangleright$ Index of linear subregion
@@
4:for $n=0:N_{max}$ do
5:  $S^{\sigma}_{\text{1}}\leftarrow\textsc{PropagateToNextRegion}(S^{\sigma}_{k})$ $\triangleright$ Propagate points to next subregion
6:  $F^{\sigma}\leftarrow\textsc{Step}(S^{\sigma}_{\text{1}})$ $\triangleright$ backward or forward step for flow vectors
7:  $\{F^{\sigma}\}_{m=0}^{M},\{S^{\sigma}_{\text{1}}\}_{m=0}^{M}\leftarrow\textsc{UniqueRegions}()$ $\triangleright$ Sort points into unique subregions
8:  $\{E^{\sigma}\}_{m=0}^{M},\{\lambda^{\sigma}\}_{m=0}^{M}\leftarrow\textsc{GetEigen}()$ $\triangleright$ Get eigenvectors/values in each region
9:  for $m=0:M$ do
10:   $k+=1$
11:   $Q^{\sigma}_{k+1}\leftarrow\textsc{GetManifold}(\{F^{\sigma}\}_{m},\{S^{\sigma}_{\text{1}}\}_{m},\{E^{\sigma}\}_{m},\{\lambda^{\sigma}\}_{m})$ $\triangleright$ Linear or curved
@@
12:   $S^{\sigma}_{k+1}\leftarrow\textsc{SamplePoints}(Q^{\sigma}_{k+1})$ $\triangleright$ Sample stable manifold
13:  end for
14:end for
15:return $(\{S^{\sigma}_{k}\}_{k=0}^{k_{max}},\{Q^{\sigma}_{k}\}_{k=0}^{k_{max}})$ $\triangleright$ Manifolds in every region
For a PL system, locally the unstable (resp. stable) manifolds of a cyclic point $\bm{p}$} are simply given by the linear subspaces spanned by the eigenvectors with eigenvalues with positive (resp. negative) real part, and hence can easily be computed in closed form from the local Jacobians $({\bm{A}}+{\bm{W}}{\bm{D}}({\bm{p}}))$. However, as soon as we cross a border into a different linear subregion, the manifold may fold and start to follow different linear dynamics. For our algorithm, we can still exploit the fact that because the dynamics in the new subregion is again linear, the corresponding piece of the manifold will either be a (hyper-)plane segment or will have curvature along one or more directions according to Eq. 4, and will be of same dimensionality as the segment in the preceding linear subregion. To determine this segment, we only need a few support vectors to correctly position it in the new subregion. Assuming $F_{\mathbf{\theta}}$ is invertible, we generate these by forward (unstable) or backward (stable) iterating the PLRNN map $F_{\mathbf{\theta}}$. Invertibility also ensures a stable manifold continues along just one branch into a preceding subregion (instead of multiple disconnected sets).Inverting Eq. 2 yields
${\bm{z}}_{t-1}=({\bm{A}}+{\bm{W}}{\bm{D}}_{t-1})^{-1}({\bm{z}}_{t}-{\bm{h}}).$ (5)
@@
A solution to this eq. needs to be self-consistent, however, i.e. the signs of the $z_{i,t-1}$ on the l.h.s. need to be consistent with the entries $d_{i,t-1}$ in ${\bm{D}}_{t-1}$ on the r.h.s. To address this, we introduce a simple heuristic: 1. Perform a backward step using the current linear subregion ${\bm{D}}_{t}$. 2. Perform a forward step using the resulting candidate solution ${\bm{z}}^{*}_{t-1}$. 3. If ${\bm{z}}_{t}==F_{\mathbf{\theta}}({\bm{z}}^{*}_{t-1})$, we are done. 4. If not, we take the candidate’s linear subregion ${\bm{D}}({\bm{z}}^{*}_{t-1})$ to attempt a new inversion. 5. If this fails, we start checking the neighboring subregions by iteratively flipping bits in ${\bm{D}}_{t-1}$. See Algorithm 2 in Appx.H.2 for details. Generally, this algorithm works for any ReLU based RNN. Its efficiency comes from the fact that usually the candidate will lie in the correct linear subregion, and – if this is not the case – it usually crosses only into neighboring regions.
@@
Figure 1: Illustration of the iterative procedure for computing stable manifolds with subregion boundaries indicated in purple-dashed. Step 1: The stable manifold (black) computation is initialized using the stable eigenvector (blue) of the saddle point (green), and sample points (orange) are placed along it. Step 2: These points are propagated until they enter another linear subregion, where the flow field is evaluated. Step 3: Using this, a new segment of the manifold is determined. Step 4: Repeating this process iteratively reconstructs the full global structure of the stable manifold. An example for a trained model can be found in Fig. 18.
We begin by using SCYFI (Eisenmann et al., 2023) to identify all fixed and cyclic points. In the linear subregion which harbors the saddle point of interest, the stable or unstable manifold is spanned by the $n_{man}$ corresponding eigenvectors. To compute the global manifold, we sample seed points on this local manifold and propagate them backward or forward in time until a new linear subregion is reached, then take one additional step to capture the local flow. If the number of stable eigenvectors in the new subregion is the same as in the original subregion and the eigenvalues are non-degenerate, we can usually represent the hyperplane or the curved manifold using the eigenvectors (Eq. 4). Otherwise we perform PCA (or kernel-PCA in the curved case), using the first $n_{man}$ components to span the manifold. We repeat this process, resampling and propagating, until all subregions have been visited or a desired depth is achieved (see Algo. 1 and Fig. 1). In systems with highly disparate timescales, this sampling process progresses fast along fast eigendirections and thereby tends to underrepresent slow eigendirections, causing numerical issues and poor approximations. To address this, we adjust sampling density inversely to eigenvalue magnitude. This balances contributions from different directions, improves accuracy for the slower directions, and ensures a more uniform representation.
@@ 3.4 Enforcing map invertibility by regularization
Figure 2: A) Runtime (top-left) and reconstruction quality (bottom) as a function of latent space dimensionality $M$ for different proportions $|\mathcal{S}_{\text{reg}}|/|\mathcal{S}|$ of subregions for which invertibility was enforced by regularization ($\lambda=0.1\exp({M})$ in Eq. 6). Means across $100$ different training runs on the Lorenz-63 system $\pm$ SD are shown. Reconstruction quality was assessed through (dis-)agreement in attractor geometry (bottom-left; $D_{\text{stsp}}$, see Appx. J) and 2020-step-ahead prediction error (bottom-right). Top-right: Relative proportion of subregions with a positive determinant of the Jacobian ($\text{det}(\bm{J})>0$), $\mathcal{S}_{+}$, as a function of latent space dimensionality $M$ for different proportions $|\mathcal{S}_{\text{reg}}|/|\mathcal{S}|$. Medians $\pm$ interquartile range are shown. B) Top: 2020-step-ahead prediction error, PE(20), as a function of the number of training epochs when the invertibility regularization, Eq. 6, was turned off (blue) vs. on (orange), for a damped nonlinear oscillator. Bottom: Convergence to a predefined performance criterion ($\text{PE}(20)\leq 0.5$) was significantly faster with the regularization turned on vs. off. Median across 2020 trained models, error bands = interquartile range. C) Same as B for Lorenz-63 system.
@@
computed across a small subset $\mathcal{S_{\text{reg}}}$ of linear subregions, where ${\bm{J}}_{i}={\bm{A}}+{\bm{W}}{\bm{D}}_{i}$ is the Jacobian in subregion $i$, and $\lambda$ a regularization parameter. As shown in Fig. 2A, strategic sampling of only $1\%$ of linear subregions, which are traversed by actual trajectories (cf. (Brenner et al., 2024a)), hardly affects runtime (Fig. 2A, top) and reconstruction performance (Fig. 2A, bottom) while still ensuring almost full invertibility (Fig. 2A, right). Since invertible flows are an inherent property of many, if not most, dynamical systems of scientific interest, we would furthermore expect that this regularization does not hamper, or even improves, the reconstruction of dynamical systems from data, in particular if the systems carry an intrinsic time reversibility (Huh et al., 2020). This is confirmed in Fig. 2B-C which shows that with the invertibility regularization in place, training of a PLRNN on a 1010d damped nonlinear oscillator converges significantly faster to a good solution ($p=0.023$, Mann-Whitney U-test) than without the regularization (Fig. 2B), while hardly affecting performance on other systems like the chaotic Lorenz-63 (Lorenz, 1963) system (Fig. 2C; $p=0.122$).
@@ 4 Delineating basins of attraction
We first validate our algorithm on a simple toy example, a two-dimensional PL map for which we have analytical forms for the inverse and the fixed points (see Appx. D.1 for details). We chose parameters to produce a simple test case with only a single saddle point. Fig. 3A confirms that points produced by backward (resp. forward) iterating the map all lie on the stable (resp. unstable) manifold as determined by our algorithm. In this simple 22d example, the manifolds correspond to line segments. Changing the PL map’s parameters slightly, we obtain a stable fixed point coexisting with a saddle period-4 (Fig. 3B, left) or period-3 (Fig. 3B, right) cyclic point (Gardini et al., 2009). Tracing back the stable manifold of the period-4 or period-3 saddle using our algorithm, we obtain the boundaries of the basins of attraction of the resp. fixed point (Fig. 3B). Fig. 3B further confirms this solution for the basin perfectly agrees with the ‘classical’ (not scalable) numerical approach of drawing initial conditions on a grid in state space, and observing their behavior in the limit $t\to\infty$.
@@
The Duffing (1918) system is a simple 22d nonlinear oscillator that can exhibit bistability between two spiral point attractors in certain parameter regimes (see Appx. I.1 for more details). We trained a shPLRNN ($M=2$, $H=10$; (Hess et al., 2023)) on this system using sparse teacher forcing (Mikhaeil et al., 2022), and used SCYFI (Eisenmann et al., 2023) to determine the fixed points. The basin boundary between the two spiral point attractors is the stable manifold of a saddle node in the center (Fig. 4A, and – as computed by our algorithm – agrees with the trajectory flows of the true system in blue.
@@
Simple models of decision making in the brain assume multistability between several choice-specific attractor states, to which the system’s state is driven as one or the other choice materializes (Wang, 2002). We trained an ALRNN ($M=15$, $P=6$) (Brenner et al., 2024a) to perform a simple 2-choice decision making task taken from (Gerstner et al., 2014), and, as before, use SCYFI to find fixed points and algorithm 1 to determine the stable manifold of a saddle separating the two basins of attraction corresponding to the two choices. The basin boundary consists of different linear pieces (hyperplanes) and is visualized in Fig. 4B in a 33d subspace (of the 1515d system) together with the reconstructed system’s fixed points (green) and some trajectories of the true system in blue.
@@
The Lorenz-63 model of atmospheric convection is probably the most famous example of a chaotic attractor. We reconstruct this system with a shPLRNN ($M=3$, $H=20$). Besides the chaotic attractor and two unstable spiral points, the system has a saddle node for which we compute the stable and unstable manifolds. Fig. 4C confirms that the manifolds computed by algorithm 1 for the shPLRNN (left), agree well with those determined by numerical integration of the original Lorenz ODE system (right). This example also proves that the shPLRNN faithfully captured the geometrical structure of the state space, beyond just reconstruction of the chaotic attractor itself.
@@
In Figure 4D we trained an ALRNN ($M=25$, $P=6$) on membrane potential recordings from a cortical neuron (Hertäg et al., 2012). The trained ALRNN contains a 38-cycle which corresponds to the rhythmic spiking activity in the real cell. In addition it has a stable fixed point and a saddle whose stable manifold (determined by our algorithm) separates the stable cycle from the stable fixed point, as illustrated in Fig. 4D (right), where we visualized the manifold through a locally linear approximation. Although we can compute the full global $24$-dimensional manifold, its inherent curvature in the $25$ dimensional state space makes it impossible to visualize it directly. Many types of cortical cells exhibit this type of bistability between spiking activity and a stable equilibrium near the resting or a more depolarized potential (Izhikevich, 2007; Durstewitz and Gabriel, 2007), and this example demonstrates how our algorithm can be utilized to reveal the structure of the state space supporting this type of dynamics from real cells.

### v3

@@ 3.3 Locating un-/stable manifolds
Hyperparameters: $N_{s}$: Number of support points to be sampled
         $N_{iter}$: Maximum number of iterations
Input: $\bm{p}:$ Periodic point
    $\sigma\in\{-1,1\}$: flag indicating whether stable (+1) or unstable (-1) manifold is to be computed
    $\bm{\theta}$: PLRNN parameters
Output: $\{Q^{\sigma}\}$: Support & eigenvectors parameterizing the manifold in each subregion
@@
[…]
4:while $n\leq N_{iter}\wedge\neg\text{isempty}(K_{\text{seed}})$ do
5:  for $k\in K_{\text{seed}}$ do $\triangleright$ Loop over regions to expand from
@@
6:   $S^{\sigma}\leftarrow\textsc{SamplePoints}(Q^{\sigma}_{k},{N_{s}})$ $\triangleright$ Sample $N_{s}$ points on the manifold
7:   $\tilde{S}^{\sigma}\leftarrow\textsc{PropagateToNextRegion}(S^{\sigma},\bm{\theta})$ $\triangleright$ Propagate forward/backward using $F^{\sigma}_{\bm{\theta}}$
8:   $K_{\text{con}(k)}\leftarrow\textsc{GetSubregionID}(\tilde{S}^{\sigma})\setminus K_{\text{visited}}$ $\triangleright$ Connecting regions not yet visited
9:   for $j\in K_{\text{con}(k)}$ do $\triangleright$ Loop over connecting regions
10:     $\lambda_{j}\leftarrow\textsc{GetEigenVal}(\bm{\theta},{\bm{D}}_{j})$ $\triangleright$ Compute eigenvalues
11:     if $\text{Im}(\lambda_{j})\neq 0\vee(\text{multiplicity}_{geo}(\lambda_{j})<\text{multiplicity}_{alg}(\lambda_{j}))$ then
12:      $Q^{\sigma}\leftarrow Q^{\sigma}\cup\{\textsc{kPCA}(\tilde{S}^{\sigma}_{j}),(\tilde{S}^{\sigma}_{j})_{1}\}$ $\triangleright$ Determine manifold by kernel-PCA
13:     else
14:      $Q^{\sigma}\leftarrow Q^{\sigma}\cup\{\textsc{PCA}(\tilde{S}^{\sigma}_{j}),(\tilde{S}^{\sigma}_{j})_{1}\}$ $\triangleright$ Determine manifold by PCA
15:     end if
16:   end for
17:   $K_{\text{visited}}\leftarrow K_{\text{visited}}\cup K_{\text{con}(k)}$ $\triangleright$ Update set of visited regions
18:  end for
19:  $K_{\text{seed}}\leftarrow K_{\text{con}}$ $\triangleright$ Make all connecting regions new seed
20:end while
21:return $\{Q^{\sigma}\}$ $\triangleright$ Manifold segments in all subregions
For our manifold detection algorithm, we exploit the PL structure of PLRNNs, which allows for an analytical construction of the manifolds within each linear subregion. We further assume that $F_{\mathbf{\theta}}$ is invertible (see above), such that the global un-/stable manifolds will be connected sets. We start by using a previously proposed algorithm, SCYFI (eisenmann_bifurcations_2023, see also pals2024inferring), to locate saddle (cyclic) points $\bm{p}$ of interest, for which we would like to construct the un-/stable manifolds. We then proceed as follows (Algo. 1; see Fig. 1 for illustration of these steps, in $1d$ for clarity):
  1. 1.
    Locally, in the linear subregion harboring $\bm{p}$, the unstable (resp. stable) manifold of $\bm{p}$} is simply given by the affine subspaces spanned by the eigenvectors with absolute eigenvalues within (stable) or outside of (unstable) the unit circle, and hence can easily be computed in closed form from the local Jacobian ${\bm{A}}+{\bm{W}}{\bm{D}}({\bm{p}})$. Note this remains true for complex eigenvalues (associated with spiral points), as these always come as conjugate pairs and hence always span a planar subspace.
  2. 2.
@@
    To extend the un-/stable manifolds across the boundaries into neighboring linear regions, we randomly sample $N_{s}$ points on the manifold and propagate them forward by $F_{\mathbf{\theta}}$ (unstable) or backward by $F^{-1}_{\mathbf{\theta}}$ (stable) into neighboring regions. The hyper-parameter $N_{s}$ is chosen such that with high likelihood all neighboring subregions into which the manifold extends are reached, which is mostly fulfilled if we choose $N_{s}=k_{\text{neigh}}\times d$, the total number of neighboring subregions times the manifold dimension. [^1] If unsure, one could increase $N_{s}$ stepwise until the number of subregions reached plateaus. For the inversion of $F_{\mathbf{\theta}}$, specifically, we need to solve ${\bm{z}}_{t-1}=({\bm{A}}+{\bm{W}}{\bm{D}}_{t-1})^{-1}({\bm{z}}_{t}-{\bm{h}})$. A solution to this eq. needs to be self-consistent, i.e. the signs of the $z_{i,t-1}$ on the l.h.s. need to be consistent with the entries $d_{i,t-1}$ in ${\bm{D}}_{t-1}$ on the r.h.s. To address this, we introduce a simple heuristic: 1) Perform a backward step using the current linear subregion ${\bm{D}}_{t}$; 2) perform a forward step using the resulting candidate solution ${\bm{z}}^{*}_{t-1}$; 3) if ${\bm{z}}_{t}==F_{\mathbf{\theta}}({\bm{z}}^{*}_{t-1})$, we are done; 4) if not, we take the candidate’s linear subregion ${\bm{D}}({\bm{z}}^{*}_{t-1})$ to attempt a new inversion; 5) if this fails, we start checking the neighboring regions by iteratively flipping bits in ${\bm{D}}_{t-1}$ (see Appx. E, Algorithm 2, for details). The efficiency of this procedure comes from the fact that usually the candidate will lie in the correct linear subregion, and – if this is not the case – it usually crosses only into neighboring regions.
  3. 3.
    The manifolds cannot change their dimensionality across subregions, but they may become curved along one or more directions as they cross into a neighboring subregion (for instance, if the dynamics in the new subregion is determined by a spiral, see example in Fig. 11). In theory, we only need to compute as many points in each subregion as required to uniquely anchor the manifold segment within that region. As we know the type of curvature qua Eq. 4, ideally these would just be $d+1$ points for a $d<M$-dimensional manifold even in the curved case, see Appx. D. In practice, numerical issues (see below) and the type of approximation used may require sampling more points. If the manifold segment is non-curved, we use PCA as a quick and simple means to obtain the affine subspace from $d+1$ support points. If the manifold is curved, we use kernel-PCA (see Appx.C). The $\geq d$ eigenvectors with non-zero eigenvalues, together with at least one support point, exactly define manifold segment $Q^{\sigma}_{k}$ in subregion kk. [^2] In Appx. D we provide another approximation for the manifold segment in each subregion, which, however, comes with certain restrictions on the number of stable eigendirections.
  4. 4.
    We then iterate steps 1-3 for at most $N_{iter}$ times, recursively expanding the manifolds across the whole state space. The algorithm’s final output will be for all $K$ subregions the set of all eigenvectors with at least one support vector positioning each manifold segment.
Note that within each linear region the construction of the manifold is thus analytical from the $\geq d+1$ support points, while across regions we recursively assemble the whole manifold following Eq. 3. In systems with highly disparate timescales, however, numerical issues may arise as sampled support points mainly spread along the fast eigendirections with little variation along the slower eigendirections, leading to poor numerical approximation of a manifold. To balance the contributions from different directions we therefore adjust sampling density inversely to eigenvalue magnitude, with an additional hyperparameter $c$ controlling this weighting (see Appx.C).
@@
Figure 1: One-dimensional illustration of the iterative procedure for computing stable manifolds with subregion boundaries indicated in purple-dashed. Step 1: The stable manifold (black) is initialized using the stable eigenvector $E^{\{+1\}}$ (blue) of the saddle point (green), and sample points (orange) are placed along it. Step 2: These points are propagated until they enter another linear subregion. Step 3: A new segment of the manifold is determined. Step 4: Repeating this process iteratively reconstructs the full global structure of the stable manifold. For a trained model example see Fig. 8.
@@ 3.4 Enforcing map invertibility by regularization
Figure 2: A) Top-left: Relative proportion of subregions with a positive determinant of the Jacobian ($\text{det}(\bm{J})>0$), $\mathcal{S}_{+}$, as a function of latent space dimensionality $M$ for different proportions $|\mathcal{S}_{\text{reg}}|/|\mathcal{S}|$. Medians $\pm$ interquartile range are shown. Top-right: Runtime and reconstruction quality (bottom) as a function of latent space dimensionality $M$ for different proportions $|\mathcal{S}_{\text{reg}}|/|\mathcal{S}|$ of subregions for which invertibility was enforced by regularization ($\lambda=0.1\exp({M})$ in Eq. 5). Means across $100$ different training runs on the Lorenz-63 system $\pm$ SD are shown. Reconstruction quality was assessed through (dis-)agreement in attractor geometry (bottom-left; $D_{\text{stsp}}$, see Appx. F) and 2020-step-ahead prediction error (bottom-right). B) Top: 2020-step-ahead prediction error, PE(20), as a function of the number of training epochs when the invertibility regularization, Eq. 5, was turned off (blue) vs. on (orange), for a damped nonlinear oscillator. Bottom: Convergence to a predefined performance criterion ($\text{PE}(20)\leq 0.5$) was significantly faster with the regularization turned on vs. off. Median across 2020 trained models, error bands = interquartile range. C) Same as B for Lorenz-63 system.
@@
computed across a small subset $\mathcal{S_{\text{reg}}}$ of linear subregions, where ${\bm{J}}_{i}={\bm{A}}+{\bm{W}}{\bm{D}}_{i}$ is the Jacobian in subregion $i$, and $\lambda$ a regularization parameter. As shown in Fig. 2A (left), strategic sampling of only $1\%$ of linear subregions, which are traversed by actual trajectories (cf. (brenner2024alrnn)), is sufficient to ensure almost full invertibility of $F_{\mathbf{\theta}}$ across the whole state space. At the same time, it hardly affects runtime (Fig. 2A, top-right) and reconstruction performance (Fig. 2A, bottom). Since invertible flows are an inherent property of many, if not most, DS of scientific interest, we would furthermore expect that this regularization does not hamper, or even improves, the reconstruction of DS from data, in particular if the systems carry an intrinsic time reversibility (huh2020time). This is confirmed in Fig. 2B-C which shows that with the invertibility regularization in place, training of a PLRNN on a 1010d damped nonlinear oscillator converges significantly faster to a good solution ($p=0.023$, Mann-Whitney U-test) than without the regularization (Fig. 2B), while hardly affecting performance on other systems like the chaotic Lorenz-63 (lorenz1963deterministic) system (Fig. 2C; $p=0.122$).
@@ 4 Delineating basins of attraction
We first validate our algorithm on a simple toy example, a two-dimensional PL map for which we have analytical forms for the inverse and the fixed points (see Appx. H.1for details). We chose parameters to produce a simple test case with only a single saddle point. Fig. 3A confirms that points produced by backward (resp. forward) iterating the map all lie on the stable (resp. unstable) manifold as determined by our algorithm. In this simple 22d example, the manifolds correspond to line segments. Changing the PL map’s parameters slightly, we obtain a stable fixed point coexisting with a saddle period-4 (Fig. 3B, left) or period-3 (Fig. 3B, right) cyclic point (gardini2009connection). Tracing back the stable manifold of the period-4 or period-3 saddle using our algorithm, we obtain the boundaries of the basins of attraction of the resp. fixed point (Fig. 3B). Fig. 3B further confirms this solution for the basin perfectly agrees with the ‘classical’ (not scalable) numerical approach of drawing initial conditions on a grid in state space, and observing their behavior in the limit $t\to\infty$.
In these simple $2d$ test cases the quality of the manifold reconstruction can easily be verified visually. In higher dimensions, however, this will no longer be possible, and so we also introduce a simple metric to quantify the reconstruction quality. For a point $\bm{x}_{0}\in\mathcal{U}$ sampled randomly either from a neighborhood $\mathcal{U}$ containing the relevant data range (see Appx. G for details), or from the reconstructed manifold of a saddle point $\bm{p}$, $\bm{x}_{0}\in W^{\sigma}(\bm{p})\cap\mathcal{U}$, with $\sigma\in\{-1,+1\}$ indicating whether we are considering the unstable or stable manifold, we compute
$\delta_{\sigma}(\bm{x}_{0}):=\frac{\min_{k\sigma\geq 0}\|F_{\bm{\theta}}^{k}(\bm{x}_{0})-\bm{p}\|^{2}_{2}}{\|\bm{x}_{0}-\bm{p}\|^{2}_{2}}\in[0,1],$ (6)
For points randomly sampled from $\mathcal{U}$ or from $W^{\sigma}(\bm{p})\cap\mathcal{U}$ for the example above, Fig.6A in Appx.G shows the distribution of $\delta_{\sigma}$, confirming $\delta_{\sigma}\approx 0$ for points sampled on the manifolds, while off the manifold $\delta_{\sigma}$ naturally spreads across a broader range. For the latter we assign an index $I_{\mathcal{U}}[\delta_{\sigma}(\bm{x}_{0}\in\mathcal{U})>\delta_{\sigma}^{\text{max}}]\in\{0,1\}$ with $\delta_{\sigma}^{\text{max}}:=\max[\delta_{\sigma}(\bm{x}_{0}\in W^{\sigma}(\bm{p})\cap\mathcal{U})]$. In the following, we report $\Delta_{\sigma}:=\langle I_{\mathcal{U}}\rangle-\tilde{\delta}_{\sigma}(\bm{x}_{0}\in W^{\sigma}(\bm{p})\cap\mathcal{U})$ as quality statistic, where $\langle\cdot\rangle$ is the mean and $\tilde{\delta}_{\sigma}$ the median.
@@
The duffing1918 system is a simple 22d nonlinear oscillator that can exhibit bistability between two spiral point attractors in certain parameter regimes (see Appx. H.3 for more details). We trained a shPLRNN ($M=2$, $H=10$; hess) on this system using sparse teacher forcing (mikhaeil2021difficulty), and used SCYFI (eisenmann_bifurcations_2023) to determine the fixed points. The basin boundary between the two spiral point attractors is the stable manifold of a saddle node in the center (Fig. 4A), and – as computed by our algorithm – agrees with the trajectory flows of the true system in blue, as further confirmed by $\Delta_{\sigma}\approx 0.97$ (Appx. Table 2 & Fig.6B).
@@
Simple models of decision making in the brain assume multistability between several choice-specific attractor states, to which the system’s state is driven as one or the other choice materializes (wang2002probabilistic). We trained an ALRNN ($M=15$, $P=6$) (brenner2024alrnn) to perform a simple 2-choice decision making task taken from (gerstner2014neuronal), and, as before, use SCYFI to find fixed points and Algorithm 1 to determine the stable manifold of a saddle separating the two basins of attraction corresponding to the two choices. The basin boundary consists of different planar pieces and is visualized in Fig. 4B in a 33d subspace (of the 1515d system) together with the reconstructed system’s fixed points (green) and some trajectories of the true system in blue ($\Delta_{\sigma}\approx 0.95$ in this case, Appx. Table 2 & Fig.6C). While visualization becomes tricky in these higher-dimensional cases, the construction in Appx. D lets us compute distances to the manifold: In this example we find, for instance, that the stable nodes are about equally far from the basin boundary ($d\approx 0.39$ vs. $d\approx 0.37$), giving the two possible choices about equal weight.
@@
The Lorenz-63 model of atmospheric convection is probably the most famous example of a chaotic attractor. We reconstruct this system with a shPLRNN ($M=3$, $H=20$). Besides the chaotic attractor (blue) and two unstable spiral points (green), the system has a saddle node for which we compute the stable and unstable manifolds $(\Delta_{\sigma}\approx 0.78$, Appx. Table 2 & Fig.6D). Fig. 4C confirms that the manifolds computed by Algorithm 1 for the trained shPLRNN (left), agree even well with those determined by numerical integration of the original Lorenz ODE system (right). This example also proves that the shPLRNN faithfully captured the geometrical structure of the state space, beyond just reconstruction of the chaotic attractor itself.
@@
In Fig. 4D we trained an ALRNN ($M=25$, $P=6$) on membrane potential recordings from a cortical neuron (hertag2012approximation). The trained ALRNN contains a 38-cycle which corresponds to the rhythmic spiking activity in the real cell. In addition it has a stable fixed point and a saddle whose stable manifold (determined by our algorithm, $\Delta_{\sigma}\approx 0.76$) separates the stable cycle from the stable fixed point, as illustrated in Fig. 4D (right), where we visualized the manifold through a locally linear approximation. Although we can compute the full global $24$-dimensional manifold, its inherent curvature in the $25$ dimensional state space makes it impossible to visualize it directly. We can, however, utilize the construction in Appx. D to compute distances to the manifold, and find that the oscillation is relatively sensitive to small perturbations at its lowest point, with a minimum distance of $d\approx 3.2$ to the basin boundary, compared to $d\approx 5.4$ for the fixed point Many types of cortical cells exhibit this type of bistability between spiking activity and a stable equilibrium near the resting or a more depolarized potential (izhikevich2007dynamical; durstewitz2007dynamical), and this example demonstrates how our algorithm can be utilized to reveal the structure of the state space supporting this type of dynamics from real cells.